OpenStudy (anonymous):

determine whether f and g are inverses of each other: f(x)=3x g(x)=x/3 f(g(x))=? g(f(x))= ?

OpenStudy (raden):

f(g(x))=f(x/3)=3(x/3)=x g(f(x))=g(3x)=3x/3=x